# VisText-Mosquito: A Multimodal Dataset and Benchmark for AI-Based Mosquito Breeding Site Detection and Reasoning

### v1

@@ Abstract
Mosquito-borne diseases pose a major global health risk, requiring early detection and proactive control of breeding sites to prevent outbreaks. In this paper, we present VisText-Mosquito, a multimodal dataset that integrates visual and textual data to support automated detection, segmentation, and reasoning for mosquito breeding site analysis. The dataset includes 1,828 annotated images for object detection, 142 images for water surface segmentation, and natural language reasoning texts linked to each image. The YOLOv9s model achieves the highest precision of 0.92926 and mAP@50 of 0.92891 for object detection, while YOLOv11n-Seg reaches a segmentation precision of 0.91587 and mAP@50 of 0.79795. For reasoning generation, our fine-tuned BLIP model achieves a final loss of 0.0028, with a BLEU score of 54.7, BERTScore of 0.91, and ROUGE-L of 0.87. This dataset and model framework emphasize the theme “Prevention is Better than Cure”, showcasing how AI-based detection can proactively address mosquito-borne disease risks. The dataset and implementation code are publicly available at GitHub: https://github.com/adnanul-islam-jisun/VisText-Mosquito
Index Terms:
Mosquito Breeding Site, Object Detection, Segmentation, Natural Language Generation, Multimodal, Public Health
I Introduction
@@
Although artificial intelligence and computer vision have made significant strides, the task of accurately detecting mosquito breeding places and analyzing water surfaces remains challenging. Existing models often underperform due to a lack of specialized datasets capable of handling both detection and segmentation tasks. Moreover, most datasets are unimodal, lacking contextual explanations that could aid in human interpretation or decision-making [4]. This absence of interpretability limits the utility of AI models in real-world public health scenarios, where both performance and explainability are critical.
To address these gaps, we introduce VisText-Mosquito, a multimodal dataset specifically designed to support three interrelated tasks: (i) detection of mosquito breeding sites, (ii) segmentation of water surfaces within these sites, and (iii) generation of natural language reasoning that describes the visual content and justifies AI predictions. The dataset comprises three main parts:
  • Breeding Place Detection: Includes 1,828 images with 3,752 annotations across five classes: coconut_exocarp, vase, tire, drain_inlet, and bottle.
  • Water Surface Segmentation: Contains 142 images with 253 annotations for two classes: vase_with_water and tire_with_water.
  • Textual Reasoning: Each image is also paired with a human-authored natural language explanation, enabling models to learn visual-linguistic associations and produce interpretable reasoning.
This paper is structured as follows: Section II reviews existing related works. Section III describes the methodology employed in our study. Section IV presents the result analysis, discussing the performance of our models across all modalities. Finally, Section V concludes the paper with a summary of contributions and future research directions.
II Existing Works
Mosquito-borne diseases remain a significant global health threat, necessitating innovative AI-driven approaches for early detection and control. Machine learning models and high-quality datasets play a pivotal role in identifying mosquito breeding sites, enabling proactive interventions. Several datasets facilitate this research, including MosquitoFusion [5], M. Mehra et al. [6], and Chathura et al. [7], which capture diverse breeding environments and enhance model adaptability.
Object detection models like YOLOv8 have demonstrated effectiveness, achieving an mAP@50 of 57.1% on the MosquitoFusion dataset. CNN-based segmentation models have also been used to identify stagnant water surfaces in drone and satellite imagery with high precision [8]. Transformer-based approaches such as DETR and SAM further improve feature representation and segmentation performance in mosquito habitat analysis tasks [9]. Meanwhile, geospatial AI techniques leveraging satellite and UAV imagery have been successfully applied to GIS-based mosquito risk mapping, enabling detection and classification of urban breeding hotspots with spatial accuracy [10]. Additionally, multimodal fusion approaches that incorporate weather, topographical, and epidemiological data have enhanced vector prediction models for disease outbreak mapping and control [11].
Despite these advancements, notable gaps remain. Many existing datasets are limited to controlled environments and focus on single-class detection, which reduces their generalizability to complex, real-world scenarios [12]. Most critically, and to the best of our knowledge, no prior work has introduced a multimodal dataset that integrates both visual (image-based detection and segmentation) and textual (natural language reasoning) components for mosquito breeding site analysis. This represents a significant gap in the field, as combining vision and language could improve interpretability, model trust, and real-world applicability of AI-driven mosquito surveillance systems.
@@ II Existing Works
Figure 2: Fully tagged and labeled images
III Methodology
In this section, we discuss the detailed methodology of our proposed solution shown in Figure 1.
III-A Data Collection
@@
The initial dataset comprises 1,828 images with 3,752 annotations for breeding place detection and 142 images with 253 annotations for water surface segmentation. This ensures a diverse, comprehensive foundation for training models on mosquito habitat detection and surface segmentation. Additionally, a text modality is included in the dataset to enable multimodal analysis. This dataset contains 3,762 instances, each associated with an image and annotated with three text fields: (a) Question: A binary question asking whether the image shows a mosquito breeding site. (b) Response: A ’Yes’ or ’No’ answer (3,748 “Yes” and 14 “No” responses). (c) Reasoning: A short free-text explanation justifying the response. The average length of the reasoning statements is approximately 36 words. This textual annotation provides semantic context and interpretability, significantly enhancing the dataset’s capacity for explainable AI.
III-B Data Preprocessing
@@
III-C Distribution Analysis and Folder Structure
The breeding place detection subset of the dataset comprises a total of 1,828 images with 3,752 annotations distributed across five classes. The class-wise distribution indicates that the Coconut-Exocarp class has the highest number of instances with 923 annotations, followed closely by the Vase class with 911 annotations. The Tire class contains 780 annotations, while the Drain-Inlet and Bottle classes have 585 and 553 annotations, respectively. For the segmentation part of the dataset, there are 142 images with a total of 253 annotations across two classes: vase_with_water and tire_with_water. The vase_with_water class has a significantly higher number of annotations, with 181 instances, compared to the tire_with_water class, which contains 72 annotations. Table I summarizes the class-wise annotation distribution in our dataset.
TABLE I: Annotation Distribution in VisText-Mosquito Dataset
Category Class Annotations
@@ III-C Distribution Analysis and Folder Structure
In addition to visual data, the dataset contains textual annotations in the form of reasoning responses that describe the rationale behind each detection or segmentation. Analysis of the reasoning texts reveals that the average length is approximately 230 characters, with most entries ranging between 175 and 280 characters. The text lengths follow a roughly normal distribution, indicating consistency in the annotation style. Most frequently occurring terms in the reasoning responses include phrases such as “stagnant water,” “mosquitoes,” “coconut shell,” and “potential breeding site,” reflecting common descriptors and domain-specific language used during the annotation process.
@@
  • images: This folder houses the visual data in the form of images collected from various mosquito breeding sites.
@@
III-D Experimental Setup
All the images in the dataset are manually reviewed to ensure that no individually identifiable information was included or embedded in the dataset. This careful review process is implemented to maintain privacy standards and ensure the dataset’s suitability for training deep learning models.
For the breeding place detection part, object detection models are trained using pre-trained versions of the YOLOv5s, YOLOv8n, and YOLOv9s models. For the segmentation part, the YOLOv8x-Seg and YOLOv11n-Seg models are employed due to their advanced capabilities in pixel-level segmentation tasks, which are critical for accurately identifying water surfaces in potential mosquito breeding sites.
For the textual reasoning component, we utilize the BLIP (Bootstrapped Language Image Pretraining) [14] model to generate natural language descriptions that explain the visual content of each annotated image. It is a vision-language model (VLM) that integrates image and text understanding and is pretrained on large-scale image-caption datasets using contrastive and generative objectives. In our method, we fine-tune the BLIP model on our curated set of reasoning texts aligned with the annotated images. During training, the model learned to associate specific visual patterns, such as objects like tires or vases containing water, with semantically rich textual descriptions that reflect potential mosquito breeding risks.
The dataset is randomly split into three subsets: 70% training images, 20% validation images, and 10% test images. This split is done to ensure a comprehensive evaluation of the models’ performance, providing ample data for both training and validation while reserving a portion for unbiased testing. The training process is conducted on a Windows 11 (Version 23H2) machine equipped with the following hardware: Nvidia RTX 3070Ti GPU with 8GB of video memory and AMD Ryzen 5800X processor. The training is run for a total of 100 epochs with the input image size set to 640 pixels, and standard hyperparameters are utilized throughout the training sessions to ensure consistent and reproducible results.
IV Result Analysis
@@
In this section, we present a comprehensive performance analysis of the models used for three core tasks: object detection, water surface segmentation, and multimodal reasoning. Evaluation metrics include Precision, Recall, and Mean Average Precision at 50% Intersection over Union (mAP@50), commonly adopted for assessing object detection and segmentation, alongside BLEU, BERTScore, ROUGE-L, and final loss for text generation evaluation.
IV-A Object Detection Performance
@@
To detect potential mosquito breeding containers, we train and evaluate three object detection models, YOLOv5s, YOLOv8n, and YOLOv9s, on 1,828 annotated images across five object classes. The detailed performance is summarized in Table II.
TABLE II: Object Detection Model Performance
@@
The YOLOv9s model achieves the highest precision and mAP@50, demonstrating its superior ability to accurately localize and classify breeding-related objects. In contrast, YOLOv5s offers the most balanced performance, maintaining high recall and demonstrating stable predictions across all classes, making it suitable for applications where minimizing false negatives is essential. YOLOv8n performed slightly lower in all metrics, suggesting that the architectural advancements in YOLOv9s provide a meaningful edge in complex real-world imagery.
IV-B Water Surface Segmentation Performance
The segmentation task is focused on identifying water surfaces in objects like vases and tires, critical indicators of mosquito breeding potential. We evaluate two advanced models, YOLOv8x-Seg and YOLOv11n-Seg, on 142 images annotated for vase_with_water and tire_with_water. Table III presents their performance:
TABLE III: Segmentation Model Performance
@@ IV-B Water Surface Segmentation Performance
YOLOv11n-Seg consistently outperform YOLOv8x-Seg across all three metrics. The improved recall indicates that YOLOv11n-Seg was more effective in correctly segmenting water surfaces without missing positive cases, which is essential in public health applications. The marginal gains in mAP@50 suggest greater consistency in its pixel-level predictions, which is especially important in distinguishing water patches under occlusion, varying illumination, or cluttered backgrounds.
IV-C Multimodal Reasoning Performance
In the textual reasoning task, we fine-tune the BLIP model to generate natural language justifications for each detection. This model is trained on image-reasoning pairs from the dataset to map visual patterns to semantically aligned textual outputs. After training, the model achieves a final loss of 0.0028, indicating successful convergence. The quality of generated reasoning is quantitatively evaluated using BLEU, BERTScore, and ROUGE-L metrics, as shown in Table IV.
TABLE IV: Multimodal Reasoning Performance (BLIP Model)
Metric Score
Loss (Final) 0.0028
BLEU 54.7
BERTScore 0.91
ROUGE-L 0.87
The high BERTScore (0.91) and ROUGE-L (0.87) indicate that the generated reasoning texts closely matched the semantic and structural properties of the ground truth. The BLEU score of 54.7 confirms strong n-gram overlap, which is valuable in capturing factual consistency. Qualitative reviews of sample outputs showed that BLIP could correctly contextualize key visual cues, such as stagnant water, container types, and environmental clutter, into meaningful, human-readable explanations, enhancing model transparency.
The combined results from detection, segmentation, and reasoning tasks validate the robustness and applicability of the VisText-Mosquito dataset and the supporting models. The YOLOv9s and YOLOv11n-Seg models exhibited state-of-the-art performance in their respective tasks, while the BLIP model added interpretability by providing contextual descriptions aligned with public health goals. This multimodal pipeline offers a powerful toolset for scalable mosquito surveillance and control systems, bridging the gap between detection accuracy and decision-making transparency.
V Conclusion
In conclusion, the VisText-Mosquito dataset provides a comprehensive multimodal resource for advancing AI-driven mosquito breeding site detection, water surface segmentation, and natural language reasoning. YOLOv9s achieve the highest precision and mAP@50 for breeding place detection, while YOLOv11n-Seg delivers superior segmentation performance. Additionally, the fine-tuned BLIP model effectively generated human-like reasoning, demonstrating the feasibility of integrating visual and textual modalities in public health. These findings emphasize the value of multimodal learning in enabling early and accurate identification of mosquito breeding sites. In the future, we plan to expand the dataset across more ecological settings and object categories, improve the diversity and quality of reasoning annotations, and explore prompt-based large language models for generating region-specific intervention strategies.

### v2

# VisText-Mosquito: A Unified Multimodal Benchmark Dataset for Visual Detection, Segmentation, and Textual Reasoning on Mosquito Breeding Sites

@@ Abstract
Mosquito-borne diseases pose a major global health risk, requiring early detection and proactive control of breeding sites to prevent outbreaks. In this paper, we present VisText-Mosquito, a multimodal dataset that integrates visual and textual data to support automated detection, segmentation, and reasoning for mosquito breeding site analysis. The dataset includes 1,828 annotated images for object detection, 142 images for water surface segmentation, and natural language reasoning texts linked to each image. The YOLOv9s model achieves the highest precision of 0.92926 and mAP@50 of 0.92891 for object detection, while YOLOv11n-Seg reaches a segmentation precision of 0.91587 and mAP@50 of 0.79795. For reasoning generation, we tested a range of large vision-language models (LVLMs) in both zero-shot and few-shot settings. Our fine-tuned Mosquito-LLaMA3-8B model achieved the best results, with a final loss of 0.0028, a BLEU score of 54.7, BERTScore of 0.91, and ROUGE-L of 0.85. This dataset and model framework emphasize the theme “Prevention is Better than Cure”, showcasing how AI-based detection can proactively address mosquito-borne disease risks. The dataset and implementation code are publicly available at GitHub: https://github.com/adnanul-islam-jisun/VisText-Mosquito
1 Introduction
@@
Although computer vision have made significant strides, the task of accurately detecting mosquito breeding places and analyzing water surfaces remains challenging. Existing models often underperform due to a lack of specialized datasets capable of handling both detection and segmentation tasks[17]. Moreover, most datasets are unimodal, lacking contextual explanations that could aid in human interpretation or decision-making [9]. This absence of interpretability limits the utility of AI models in real-world public health scenarios, where both performance and explainability are critical.
To address these gaps, we present VisText-Mosquito, the first multimodal benchmark for mosquito breeding site analysis. Our dataset integrates three complementary components: (i) object detection with 1,828 images and 3,752 bounding boxes across five container types (coconut_exocarp, vase, tire, drain_inlet, and bottle); (ii) water surface segmentation with 142 high-resolution images and 253 pixel-level masks; and (iii) natural language reasoning with automatically generated and human-validated explanations that capture visual cues linked to mosquito breeding risk. Together, these components enable joint evaluation of detection, segmentation, and multimodal reasoning in a unified benchmark.
Beyond dataset creation, we perform a comprehensive benchmarking study of state-of-the-art detectors (YOLOv5s/8n/9s), segmenters (YOLOv8x-Seg, YOLOv11n-Seg), and large vision–language models (LLaVA, DeepSeek-VL2, InternVL-4B, Gemini-2.5-Flash). Finally, we fine-tune LLaMA3-8B-Vision on our reasoning corpus, yielding Mosquito-LLaMA3-8B, which achieves state-of-the-art BLEU, ROUGE-L, and BERTScore, demonstrating the value of domain adaptation for multimodal reasoning. Our key contributions are as follows:
(1) We introduce VisText-Mosquito, the first multimodal dataset that unifies object detection, water segmentation, and visual–linguistic reasoning for mosquito breeding site analysis. (2) We provide high-quality annotations across visual and textual modalities, including bounding boxes, segmentation masks, and reasoning texts for explainable AI, which are all human-annotated and validated. (3) We benchmark a wide spectrum of detectors, segmenters, and VLMs under zero-shot and few-shot setups, revealing the limitations of unimodal approaches. (4) We develop Mosquito-LLaMA3-8B, a domain-adapted VLM that sets a new baseline for multimodal reasoning in public health surveillance.
2 Related Works
Mosquito-borne diseases continue to pose serious public health threats globally, especially in tropical and subtropical regions [20]. Effective surveillance of mosquito breeding habitats is essential for early intervention and disease prevention. In recent years, advances in computer vision and machine learning have enabled automated detection of mosquito habitats using images and spatial data.
Several datasets have been proposed to support AI-based mosquito breeding site detection. The MosquitoFusion dataset [17] offers annotated real-world images capturing mosquitoes, larvae, swarms, and breeding sites across diverse contexts. Mehra et al. [13] created an entomological image dataset for training lightweight object detection models under challenging lighting and occlusion conditions. Similarly, Chathura et al. [19] presented a drone-captured image dataset that includes standing water sources and vegetation typically associated with mosquito breeding, enhancing the generalizability of deep learning models to outdoor environments.
Object detection architectures such as YOLOv5, YOLOv8, and EfficientDet have shown promising results in localizing potential breeding containers, with YOLOv8 achieving a mean average precision (mAP@50) of 57.1% on the MosquitoFusion dataset [17]. For semantic and instance segmentation, CNN-based models like U-Net and DeepLabV3+ have been used to segment water surfaces from drone and satellite imagery with pixel-level accuracy [12, 15]. Recent transformer-based models like DETR [14] and the Segment Anything Model (SAM) [8] offer superior segmentation quality by learning rich visual features from complex environments.
Geospatial AI techniques have been applied to UAV and satellite imagery for spatial mapping of mosquito breeding hotspots. For instance, GIS-based models using remote sensing data and LIDAR have enabled high-resolution risk mapping of urban habitats [4, 18]. Studies have also explored topographical features, vegetation indices, and hydrological data to predict stagnant water accumulation areas likely to host larvae [7]. Moreover, ensemble machine learning methods (e.g., Random Forest, XGBoost) using environmental and epidemiological data have improved mosquito vector prediction models, especially when coupled with weather and public health surveillance data [2].
In the multimodal domain, progress has been slower. Most existing works focus solely on image-based detection or segmentation, with limited incorporation of textual or reasoning-based outputs. However, natural language generation (NLG) in vision-language tasks has gained popularity, especially with models like BLIP [10], LLaVA [11], and Flamingo [1], which generate human-like captions or justifications. These models have been applied to medical imaging, environmental monitoring, and accessibility applications, showing promise for tasks requiring explainability and human alignment.
Despite these advancements, notable limitations persist. Many datasets are restricted to lab environments or synthetic scenes, lacking the visual complexity of urban and peri-urban breeding sites. Furthermore, most existing models treat breeding site identification as a visual-only task, ignoring the value of language-based reasoning for public health decision-making and education. To the best of our knowledge, no prior work has introduced a multimodal dataset that jointly supports object detection, water surface segmentation, and visual reasoning tailored to mosquito breeding site analysis. This represents a significant research gap, as integrating vision and language could boost both interpretability and field applicability of AI models for mosquito vector control.
3 Dataset Construction Process
@@
Figure 1: Fully tagged and labeled images for breeding site detection and surface water segmentation
3.1 Data Collection
@@
The initial dataset comprises 1,828 images with 3,752 annotations for breeding place detection and 142 images with 253 annotations for water surface segmentation. This ensures a diverse, comprehensive foundation for training models on mosquito habitat detection and surface segmentation. Additionally, a text modality is included in the dataset to enable multimodal analysis. This dataset contains 3,762 instances, each associated with an image and annotated with three text fields: (a) Question: A binary question asking whether the image shows a mosquito breeding site. (b) Response: A ’Yes’ or ’No’ answer (3,748 “Yes” and 14 “No” responses). (c) Reasoning: A short free-text explanation justifying the response. The average length of the reasoning statements is approximately 36 tokens. This textual annotation provides semantic context and interpretability, significantly enhancing the dataset’s capacity for explainable AI.
3.2 Data Preprocessing
@@
3.3 Distribution Analysis and Folder Structure
The breeding place detection subset of the dataset comprises a total of 1,828 images with 3,752 annotations distributed across five classes. The class-wise distribution indicates that the Coconut-Exocarp class has the highest number of instances with 923 annotations, followed closely by the Vase class with 911 annotations. The Tire class contains 780 annotations, while the Drain-Inlet and Bottle classes have 585 and 553 annotations, respectively. For the segmentation part of the dataset, there are 142 images with a total of 253 annotations across two classes: vase_with_water and tire_with_water. The vase_with_water class has a significantly higher number of annotations, with 181 instances, compared to the tire_with_water class, which contains 72 annotations. Table 1 summarizes the class-wise annotation distribution in our dataset.
Category Class # Anno.
@@ 3.3 Distribution Analysis and Folder Structure
Table 1: Annotation Distribution in VisText-Mosquito Dataset
In addition to visual data, the dataset contains textual annotations in the form of reasoning responses that describe the rationale behind each detection or segmentation. Analysis of the reasoning texts reveals that the average length is approximately 230 characters, with most entries ranging between 175 and 280 characters. The text lengths follow a roughly normal distribution, indicating consistency in the annotation style. Most frequently occurring terms in the reasoning responses include phrases such as “stagnant water,” “mosquitoes,” “mosquito larvae,” and “potential breeding site,” reflecting common descriptors and domain-specific language used during the annotation process. A word cloud is generated to visualize the most frequently occurring terms in the reasoning responses, highlighting key descriptors and domain-specific language used during the annotation process. Figure 2 shows the word cloud derived from the textual reasoning data.
Figure 2: Word Cloud of Common Terms in Reasoning Texts
@@
  • images: This folder contains the visual data in the form of images collected from various mosquito breeding sites.
@@
4 Experimental Setup
In this section, we discuss the detailed experimental setup of our proposed solution shown in Figure 3.
4.1 Hardware Configurations
All images were manually reviewed to ensure no personally identifiable information was present. The dataset was split into 70% training, 20% validation, and 10% testing. For object detection and segmentation, models were trained for 100 epochs on a Linux workstation with Ubuntu 24.04, equipped with an NVIDIA RTX 3090 Ti GPU (24 GB) and an AMD Ryzen 5800X CPU. Images were resized to 640×640 pixels, and standard YOLO hyperparameters were used for consistency across experiments. For multimodal reasoning, we fine-tuned LLaMA3-8B-Vision for 3 epochs using image–question–reasoning triplets. Training employed the AdamW optimizer with a learning rate of $5\times 10^{-5}$ and batch size of 16.
4.2 Models
For the breeding site detection task, we use three state-of-the-art object detection models: YOLOv5s, YOLOv8n, and YOLOv9s. These models are chosen for their lightweight architecture, fast inference, and competitive accuracy, especially for detecting small, object-like containers or wet surfaces associated with mosquito breeding. In parallel, we apply YOLOv8x-Seg and YOLOv11n-Seg for the water surface segmentation task. These models extend object detection with dense pixel-wise segmentation capabilities, enabling precise boundary-level understanding of water-holding structures such as tires and vases. All models are initialized with pre-trained weights from the COCO dataset and then fine-tuned on the VisText-Mosquito dataset to adapt to our specific domain.
To support multimodal understanding, we explore a wide range of large vision-language models for the task of generating natural language reasoning based on images. These models are tasked with answering a yes/no question and generating free-form textual justifications grounded in visual evidence. We evaluate both open-source and closed-source VLMs. From the closed-source domain, we test Gemini-2.5-Flash, which is a proprietary model. For open-source alternatives, we evaluate LLaVA-1.5-7B, LLaMA3-8B-Vision, DeepSeek-VL2, InternVL-4B, and Qwen2-VL.
4.3 Prompting Strategies
Each vision-language model is tested under two prompting strategies: zero-shot and few-shot. In the zero-shot setting, the model is given only the input image and question, without any in-context examples. This setup evaluates the model’s out-of-the-box generalization capacity to unseen prompts. In the few-shot setting, we include three image-question-reasoning pairs as exemplars in the prompt to guide the model.
4.4 Supervised Fine-Tuning
Beyond prompting, we perform supervised fine-tuning using the LLaMA3-8B-Vision model on the VisText-Mosquito dataset, which is named Mosquito-LLaMA3-8B. During fine-tuning, the model learns to generate reasoning text conditioned on the image content and question prompt using a cross-entropy loss over token probabilities. The model is optimized using the AdamW optimizer with a learning rate of $5\times 10^{-5}$ and trained using a batch size of 16. Training proceeds for 3 epochs. This process enables the model to capture domain-specific visual-textual associations and produce reliable and interpretable reasoning.
4.5 Evaluation Metrics
To measure the effectiveness of our models, we adopt both vision-based and language-based evaluation metrics. For object detection and segmentation tasks, we report precision, recall, and mean Average Precision at IoU threshold 0.5 (mAP@50). These metrics provide insight into how accurately the models detect and localize breeding sites and water surfaces. For the reasoning generation task, we use a combination of BLEU, ROUGE-L, and BERTScore to evaluate the quality of generated text. BLEU captures n-gram overlap with reference reasoning, ROUGE-L assesses longest common subsequences to reflect fluency and relevance, while BERTScore leverages contextual embeddings to capture semantic similarity. Together, these metrics provide a holistic view of textual generation quality in terms of lexical, structural, and semantic alignment.
5 Result Analysis
Figure 4: Confusion matrices for (Top) object detection and (Bottom) segmentation tasks.
@@ 5 Result Analysis
In this section, we present a comprehensive performance analysis of the models used for three core tasks: object detection, water surface segmentation, and multimodal reasoning.
5.1 Object Detection Performance
@@
To detect potential mosquito breeding sites, we fine-tune and evaluate three object detection models, YOLOv5s, YOLOv8n, and YOLOv9s, on 1,828 annotated images across five object classes. The detailed performance is summarized in Table 2.
Table 2: Object Detection Model Performance
@@
The YOLOv9s model achieves the highest precision of 0.92926 and mAP@50 of 0.92891, demonstrating its superior ability to accurately localize and classify breeding-related objects. In contrast, YOLOv5s offers the most balanced performance, with a strong precision of 0.91514 and the highest recall of 0.87595. This makes it particularly well-suited for applications where minimizing false negatives is essential, such as early-stage mosquito surveillance. Meanwhile, YOLOv8n shows slightly lower performance across all metrics, with a precision of 0.89028, recall of 0.87314, and mAP@50 of 0.90817. These results suggest that the architectural improvements in YOLOv9s contribute to meaningful gains in real-world breeding site detection tasks.
5.2 Water Surface Segmentation Performance
The segmentation task is focused on identifying water surfaces in objects like vases and tires, critical indicators of mosquito breeding potential. We fine-tuned two advanced models, YOLOv8x-Seg and YOLOv11n-Seg, on 142 images annotated for vase_with_water and tire_with_water. Table 3 presents their performance:
Table 3: Segmentation Model Performance
@@ 5.2 Water Surface Segmentation Performance
Model Zero-shot Few-shot
BLEU BERTScore ROUGE-L BLEU BERTScore ROUGE-L
Gemini-2.5-Flash 45.9 0.78 0.73 47.5 0.83 0.77
LLaVA-1.5-7B 38.2 0.66 0.65 39.1 0.68 0.69
LLaMA3-8B-Vision 40.2 0.71 0.68 42.8 0.78 0.72
DeepSeek-VL2 36.7 0.64 0.63 38.2 0.73 0.66
InternVL-4B 33.5 0.52 0.57 35.4 0.56 0.60
Qwen2-VL 31.6 0.55 0.60 29.3 0.50 0.55
Mosquito-LLaMA3-8B (Ours) 41.4 0.87 0.78 54.7 0.91 0.85
Table 4: Zero-shot and Few-shot Reasoning Performance Across Vision-Language Models. Our fine-tuned Mosquito-LLaMA3-8B achieves top scores under few-shot conditions.
YOLOv11n-Seg consistently outperforms YOLOv8x-Seg across all three evaluation metrics. Specifically, it achieves a higher precision of 0.91587 compared to 0.89372, and a notably higher recall of 0.77201 versus 0.73074, reflecting its superior ability to detect true positive water surface segments without missing relevant regions. Additionally, the improvement in mAP@50 from 0.79345 to 0.79795, although modest, suggests more stable and consistent segmentation performance, particularly beneficial in challenging real-world settings with occlusion, shadows, or visually cluttered environments. These results underline the robustness of YOLOv11n-Seg for pixel-level detection tasks in mosquito habitat identification.
5.3 Multimodal Reasoning Performance
Configuration BLEU BERTScore ROUGE-L Final Loss
Baseline (3 epochs, full data) 54.7 0.91 0.85 0.0028
1 epoch (underfitting) 46.2 0.82 0.76 0.0085
5 epochs (overfitting) 49.7 0.87 0.81 0.0016
Learning rate $1\times 10^{-5}$ 47.5 0.85 0.79 0.0043
Learning rate $1\times 10^{-4}$ 48.8 0.86 0.80 0.0039
50% training examples 42.1 0.76 0.72 0.0071
75% training examples 48.3 0.84 0.77 0.0052
No image augmentation 46.8 0.81 0.75 0.0060
Table 5: Ablation Study on Fine-Tuned Mosquito-LLaMA3-8B
For the textual reasoning task, we fine-tune the LLaMA3-8B-Vision model using image-question-reasoning triplets curated from the VisText-Mosquito dataset. The model is trained to generate coherent and semantically grounded justifications that explain whether an image depicts a mosquito breeding site and why. During training, it learns to associate domain-specific visual patterns, such as standing water in tires, vases, or drains, with meaningful language that reflects expert-level reasoning. After three epochs, the model converges with a final training loss of 0.0028, suggesting stable training and strong multimodal alignment.
To evaluate the generated justifications, we adopt standard natural language generation metrics including BLEU, BERTScore, and ROUGE-L. Our fine-tuned model, referred to as Mosquito-LLaMA3-8B, achieves a BLEU score of 54.7, a BERTScore of 0.91, and a ROUGE-L of 0.85 in the few-shot setting. These results indicate strong n-gram precision, semantic similarity, and structural alignment with human-written reference texts. Qualitative inspection further supports these findings: the model effectively integrates visual context, such as the presence of water, container types, and environmental cues, into fluent, human-readable rationales.
Beyond supervised fine-tuning, we assess the generalization and adaptation abilities of other vision-language models under both zero-shot and few-shot prompting setups. In the zero-shot setting, models receive only an image and a prompt without any additional context. In the few-shot setting, each model is shown three image-question-reasoning examples before generating responses for unseen inputs. This experimental design enables us to evaluate both generalization from pretraining and responsiveness to in-context learning.
Table 4 summarizes the results. Overall, few-shot prompting improved performance across most models, with notable increases in BLEU, BERTScore, and ROUGE-L. For instance, LLaMA3-8B-Vision improved from 40.2 to 42.8 in BLEU, and from 0.71 to 0.78 in BERTScore. LLaVA-1.5-7B and DeepSeek-VL2 showed modest but consistent improvements across all metrics, while even the smaller InternVL-4B gained from few-shot examples, despite its relatively lower baseline. Interestingly, the closed-source Gemini-2.5-Flash displayed only minor improvements, possibly due to its already strong zero-shot capabilities limiting headroom for further gain.
A notable exception to this trend is Qwen2-VL, which experienced a performance decline in all three metrics under few-shot prompting, dropping from 31.6 to 29.3 in BLEU, from 0.55 to 0.50 in BERTScore, and from 0.60 to 0.55 in ROUGE-L. This suggests that Qwen2-VL may be particularly sensitive to prompt construction or lacks sufficient in-context learning capabilities for domain-specific reasoning tasks.
In contrast, our fine-tuned Mosquito-LLaMA3-8B model exhibits significant improvements across all metrics in the few-shot setting. Compared to its own zero-shot variant, it achieves a +13.3 point gain in BLEU, +0.04 in BERTScore, and +0.07 in ROUGE-L. These results highlight the clear advantage of supervised domain-specific fine-tuning over reliance on general pretraining or prompt-based learning. Overall, our findings indicate that while few-shot prompting enhances reasoning capabilities in most models, fine-tuning remains essential for achieving high-accuracy, context-sensitive reasoning in public health surveillance applications.
5.4 Ablation Study
To understand the effect of different training configurations on reasoning performance, we conduct an ablation study on the fine-tuned Mosquito-LLaMA3-8B model. We vary critical training hyperparameters, including the number of training epochs, learning rate, and the number of image-question-reasoning triplets used during training. The goal is to isolate how each factor contributes to the model’s ability to align visual cues with textual reasoning.
Table 5 summarizes the results. The baseline model (our final version) is trained for 3 epochs with a learning rate of $5\times 10^{-5}$ and the full dataset. Reducing the number of epochs to 1 or increasing it to 5 both result in slightly degraded performance, likely due to underfitting and overfitting, respectively. Similarly, lowering the learning rate delays convergence and results in weaker semantic alignment. Using only 50% of the training examples notably degrades all metrics, demonstrating the importance of dataset size for effective domain-specific reasoning.
The ablation confirms that the final configuration strikes a strong balance between generalization and alignment. Notably, the number of training examples and moderate epoch count are the most influential factors. These findings can guide future fine-tuning efforts on similar multimodal public health datasets.
6 Discussion
Despite the strong performance of both detection and reasoning models on the VisText-Mosquito dataset, several challenges and limitations remain, particularly in the context of multimodal understanding. Through qualitative analysis, we observed recurring failure cases among VLMs, especially in zero-shot scenarios. For instance, models like Qwen2-VL and InternVL-4B often failed to generate grounded reasoning when visual cues were subtle or occluded. A common example includes misinterpreting a dry tire partially filled with leaves as a mosquito breeding site, where the model incorrectly inferred the presence of water based on contextual noise rather than actual visual evidence.
Another frequent failure pattern emerged when visually similar objects led to semantic confusion. VLMs sometimes misclassified objects such as a decorative vase or a plastic bucket as breeding containers even when they lacked water. This suggests that models are often over-reliant on coarse object shape and context rather than finer visual-textual alignment. Additionally, models struggled when backgrounds were cluttered or when key cues like water surfaces were not prominently visible. These cases highlight the limitations of current VLMs in precisely localizing fine-grained visual patterns critical to public health applications.
The disparity between zero-shot and few-shot performance across models further reinforces the importance of context-aware prompting. Closed-source models like Gemini-2.5-Flash performed reasonably well even without training, but open-source alternatives significantly improved under few-shot settings, indicating their reliance on task-specific cues. However, even in few-shot setups, hallucinations were observed, some models generated plausible yet factually incorrect reasoning, such as inferring mosquito larvae without any visible evidence.
These findings emphasize the need for further advancements in VLM robustness, particularly in environments with ambiguous visual signals. Incorporating spatial grounding, multimodal attention refinement, and post-hoc factuality checks could improve reliability. Moreover, continued dataset expansion with hard negative samples and edge-case reasoning examples may help fine-tune these models for real-world vector surveillance scenarios.
7 Conclusion
The VisText-Mosquito dataset establishes a comprehensive multimodal benchmark for automated detection, segmentation, and reasoning on mosquito breeding sites. Through rigorous experiments, we show that YOLOv9s yields the best detection accuracy, while YOLOv11n-Seg excels in water surface segmentation. Our fine-tuned Mosquito-LLaMA3-8B model shows strong performance in generating coherent justifications, bridging visual cues with textual reasoning. These results underscore the potential of integrating vision-language models into public health monitoring frameworks. By combining high-performance object recognition with explainable reasoning, our dataset paves the way for intelligent, interpretable, and scalable tools to support vector control strategies. Future work will expand the dataset across ecological regions, enhance reasoning diversity, and explore prompt-adaptive models for localized intervention planning.